# Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code

### v1

@@ Abstract
Code watermarking identifies AI-generated code by embedding patterns into the code during generation. Effective watermarking requires meeting two key conditions: the watermark should be reliably detectable, and the code should retain its original functionality. However, existing methods often modify tokens that are critical for program logic, such as keywords in conditional expressions or operators in arithmetic computations. These modifications can cause syntax errors or functional failures, limiting the practical use of watermarking. We present STONE, a method that preserves functional integrity by selectively inserting watermarks only into non-syntax tokens. By excluding tokens essential for code execution, STONE minimizes the risk of functional degradation.
In addition, we introduce CWEM, a comprehensive evaluation metric that evaluates watermarking techniques based on correctness, detectability, and naturalness. While correctness and detectability have been widely used, naturalness remains underexplored despite its importance. Unnatural patterns can reveal the presence of a watermark, making it easier for adversaries to remove. We evaluate STONE using CWEM and compare its performance with the state-of-the-art approach. The results show that STONE achieves an average improvement of 7.69% in CWEM across Python, C++, and Java. Our code is available in https://github.com/inistory/STONE-watermarking/.
[…]
linecolor=white, backgroundcolor=gray!20,
@@ 1 Introduction
The rapid development of large language models (LLMs) has significantly improved the ability to generate human-like text and code (Nam et al., 2024; Kazemitabaar et al., 2023; Liu et al., 2023a; Sun et al., 2024; Guo et al., 2024a; Wang et al., 2024; Wei et al., 2023; Tang et al., 2023; Guo et al., 2024b). These advancements have unlocked new possibilities across various fields, including software development and automated code generation. However, this progress has introduced challenges in tracing the origin of generated code (Yang et al., 2023; Li et al., 2023). As LLM-generated code becomes increasingly indistinguishable from human-written code, researchers have explored watermarking techniques to address this issue.
LLM watermarking involves embedding specific patterns into the generated output during the generation process to help identify its origin (Kirchenbauer et al., 2023; Lee et al., 2024). These patterns should be imperceptible to human reviewers but remain detectable through algorithmic analysis, enhancing the transparency and accountability of AI-generated content (Guo et al., 2024c; Hou et al., 2024).
@@
Among various approaches, the green and red list-based watermarking is the most widely adopted (Kirchenbauer et al., 2023; Zhao et al., 2023; Lee et al., 2024; Lu et al., 2024; Guo et al., 2024c). This method divides token candidates into two lists and biases the selection process toward tokens in the green list. Detection involves measuring the proportion of green list tokens in the generated output. While effective for natural language tasks, this approach does not easily apply to code generation, where structural and functional correctness are crucial (Lee et al., 2024; Hoang et al., 2024). Changes to syntax-critical tokens can cause compilation errors or alter program behavior.
Code watermarking identifies AI-generated code by embedding patterns into the code while maintaining functional correctness. Effective watermarking requires meeting two conditions: reliable watermark detectability and intact code correctness. However, many existing methods inadvertently modify tokens essential to program logic, such as keywords in conditional expressions or operators in arithmetic computations. These modifications can introduce syntax errors or change the behavior of the code, undermining its reliability. We propose STONE (Syntax TOkeN preserving codE watermarking), a method that selectively inserts watermarks into non-syntax tokens. By excluding tokens that are critical to code execution, STONE reduces the risk of functional degradation while embedding robust, detectable patterns.
Additionally, we introduce CWEM (Code Watermarking Evaluation Metric), a comprehensive framework for evaluating watermarking techniques. CWEM assesses performance across three key dimensions: correctness, detectability, and naturalness. While correctness and detectability have been the focus of prior research, naturalness is equally important. Watermarking methods that create unnatural code patterns make it easier for adversaries to detect and remove the embedded patterns. We evaluate STONE using CWEM and compare its performance with the state-of-the-art approach. Our experiments demonstrate that STONE achieves an average improvement of 7.69% in CWEM across Python, C++, and Java.
2 Preliminary analysis
@@
The previous SOTA approach SWEET (Lee et al., 2024) takes into account the structural and syntactic constraints by embedding watermarks only in high-entropy tokens, it ensures a minimal impact on the code quality. In our preliminary analysis, we group code tokens into several categories and compare their average entropy values to see if high-entropy tokens have a smaller impact on code quality. We consider two popular benchmarks, MBPP+ and HumanEval+, for the Python code generation and completion. We classify each Python code into one of five categories reported in Table 7 where any remaining tokens are put into the etc. category.
@@
where $H_{t}$ is the entropy value at generation step $t$, $V$ is the vocabulary of the language model, $y_{<t}$ is the sequence of tokens generated before step $t$, and $P(y_{t}=v_{i}\mid y_{<t})$ is the probability assigned by the model to token $v_{i}$ as the next token, respectively. The entropy $H_{t}$ measures the uncertainty of the next token prediction. A high entropy value indicates that the model is uncertain and many tokens have comparable probabilities. Conversely, low entropy implies that one token has a dominant probability.
Figure 1: Entropy values by token types. We refer to tokens other than keywords, whitespaces, types, delimiters, and operators as etc tokens.
@@
Since keywords have high entropy, embedding watermarks in high-entropy tokens, as in SWEET, carries the risk of placing watermarks on keywords. Keywords are essential to the syntax and logic of the code, thus watermarking them could negatively impact the quality and functionality of the code. In comparison, the etc category, while less critical to the functionality of the code, has relatively high entropy. Therefore, embedding watermarks in etc tokens may be a more suitable approach for preserving code quality.
Figure 2: A motivation example comparing SWEET (Lee et al., 2024) and STONE (ours). (a) An example prompt for code generation, (b) Solution code without any watermark, (c) Output code watermarked by SWEET, and (d) Output code watermarked by STONE. In this example, SWEET accidentally replaces % (modulus) with ///// / (integer division) when checking if a number is even, leading to incorrect logic.
3 Approach: STONE
STONE watermarking is a robust watermarking approach designed for preserving code quality illustrated in Figure 2. This section introduces the watermark insertion and detection process of our proposed method. In Section 3.1, we detail how watermarks are inserted by selectively modifying non-syntactic tokens while keeping critical elements like keywords, operators, and delimiters unchanged to preserve functionality and structure. In Section 3.2, we explain how the watermark is detected by analyzing the green token distribution among non-syntax elements and computing a statistical $z$-score to confirm the presence of the watermark.
3.1 Inserting Watermarks
Algorithm 1 Generation Algorithm of STONE
1:Tokenized prompt $\mathbf{x}$, previously generated tokens $\mathbf{y}_{:t}=(y_{0},\dots,y_{t-1})$, vocabulary set $\mathcal{V}$, language model $f_{\text{LM}}$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, logit adjustment factor $\delta>0$
@@ 3.1 Inserting Watermarks
7:         Compute a hash of token $y_{t-1}$
8:         Randomly divide $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
@@
We define the keywords, whitespaces, types, delimiters, and operators presented in Table 7 as functionally important syntactic elements and embed watermarks only in tokens that do not belong to these categories. Algorithm 1 describes how STONE embeds watermarks during code generation. Given a tokenized prompt $\mathbf{x}$ and $\mathbf{y}_{:t}$, the tokens generated up to the previous time step, the language model $f_{\text{LM}}$ computes the initial probability distribution $\mathbf{p}_{t}$ for the token to be generated at the current time step $t$. STONE samples a candidate token $\tilde{y}_{t}$ based on this initial probability distribution. If the sampled token is not in the syntax element set, STONE divides the vocabulary $\mathcal{V}$ into a green list and a red list according to the green token ratio $\gamma$. It then increases the likelihood of sampling tokens from the green list by adding a constant $\delta$ to their logit values. Using the adjusted logits, STONE recalculates the probability distribution and samples the final token $y_{t}$ for time step $t$. Since STONE embeds watermarks while preserving functionally important tokens, it is able to maintain the quality of the code.
3.2 Detecting Watermarks
@@
7:         Split vocabulary $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
@@
9:              $N^{E}_{G}\leftarrow N^{E}_{G}+1$triangleright Token matches the green list
@@
We assess how many green tokens are present in the code, and if this value surpasses a certain threshold, we conclude that the code was generated by an LLM. When identifying potential green tokens, we focus only on non-syntactic elements (i.e., etc tokens). Algorithm 2 details the watermark detection process for STONE scenario. $N^{E}$ represents the number of etc tokens in the code, and $N^{E}_{G}$ denotes the number of green tokens among them. We replicate the division of the vocabulary into green and red lists at each time step during code generation to determine whether a specific token is a green token. If the $z$-score exceeds the predefined threshold $z_{\text{threshold}}$, the code is determined to be generated by an LLM. The purpose of embedding and detecting watermarks is to determine whether a given piece of code was generated by an LLM.
4 Code Watermarking Evaluation Metric
Existing code watermarking research lacks a unified evaluation metric that simultaneously considers functional correctness, detectability and naturalness (Lee et al., 2024; Yang et al., 2024; Li et al., 2024; Guan et al., 2024). They have typically focused one or two of these properties, leading to incomplete assessments and conflicting interpretations. By integrating functional correctness, detectability, and naturalness, we propose CWEM (Code Watermarking Evaluation Metric), a flexible metric that balances each factor using weights $\alpha$,$\beta$ and $\gamma$ satisfying $\alpha$ + $\beta$ + $\gamma$ = 1. In practice, these coefficients reflect the relative importance of functional correctness ($\alpha$ ), watermark detectability ($\beta$) and fluency or naturalness ($\gamma$). For instance, systems that cannot risk functional errors can assign a larger $\alpha$, while environments prioritizing watermark detection may emphasize $\beta$.
[…]
$+\gamma\cdot\mathrm{Naturalness}(C_{wm},C),$
correctness is measured through the pass@$k$ change before and after watermarking (Section 4.1), detectability is quantified using $z$-score and AUROC metric (Section 4.2) and naturalness is assessed using perplexity (Section 4.3).
4.1 Correctness
Ensuring correctness is crucial to prevent watermarking from compromising the functionality of the code. We measure functionality preservation using the pass@k metric. Specifically, we adopt the unbiased version of pass@$k$ (Chen et al., 2021). In this approach, $k$ samples per problem are generated, and the number of correct samples, $c$, is counted as the number of watermarked code samples in the set ${C_{wm}}$ that pass all unit tests.
This metric estimates pass@$k$ with the formula:
@@ 4.1 Correctness
where $\mathbb{E}_{C_{wm}}$ denotes the expectation over the set of watermarked code samples $C_{wm}$, $n$ is the total number of generated code samples for a given problem, $c$ is the number of correct code samples within the set, $\binom{n}{k}$ denotes the number of ways to choose $k$ code samples from the total $n$ samples, and $\binom{n-c}{k}$ represents the number of ways to choose $k$ code samples from the incorrect samples only (i.e., excluding the $c$ correct samples). The term $1-\frac{\binom{n-c}{k}}{\binom{n}{k}}$ calculates the probability that at least one correct code sample is selected when $k$ samples are drawn from the total $n$ samples. The overall correctness metric is obtained as the expected value of this probability over the set $C_{wm}$.
4.2 Detectability
Detectability measures how effectively we can identify the presence of a watermark in generated code. We assess the detectability by computing a $z$-score based on the distribution of green-list tokens in the generated text. The classification performance is then evaluated using AUROC (Area Under the Receiver Operating Characteristic Curve).
@@ 4.2 Detectability
$Z$-score: Green Token Ratio
For a given code token sequence $C_{wm}=\{c_{0},c_{1},...,c_{T}\}$, the $z$-score is computed as:
$z(C_{wm})=\frac{|C_{wm}|_{G}-\gamma T}{\sqrt{T\gamma(1-\gamma)}},$ (4)
where $|C_{wm}|_{G}$ is the number of green-list tokens in the sequence $C_{wm}$, $T$ is the total number of tokens in the sequence, and $\gamma$ is the expected proportion of green tokens in a non-watermarked sequence. The denominator represents the standard deviation under the binomial distribution assumption. This $z$-score quantifies the deviation of the observed green token frequency from its expected value, enabling the detection of watermarked text.
LLM-Generated Code Detection Using Green Token Ratio
We evaluate classification performance using AUROC (Area Under the Receiver Operating Characteristic Curve) , which quantifies how well we can distinguish between human-written code $C_{H}$ and watermarked code $C_{wm}$ based on their $z$-scores.
We treat $z(C_{H})$ as scores for the negative class (human-written) and $z(C_{wm})$ as scores for the positive class (watermarked). The ROC curve is obtained by varying a decision threshold $\tau$ computing the corresponding True Positive Rate (TPR) and False Positive Rate (FPR). TPR is the fraction of watermarked sequences correctly identified as watermarked.
$\text{TPR}(\tau)=\frac{\sum_{j}\mathbf{1}(z(C_{wm}^{(j)})>\tau)}{|{C_{wm}}|},$ (5)
FPR is the fraction of human-written sequences incorrectly classified as watermarked.
$\text{FPR}(\tau)=\frac{\sum_{i}\mathbf{1}(z(C_{H}^{(i)})>\tau)}{|{C_{H}}|},$ (6)
For each threshold $\tau$, a sequence is classified as watermarked if its computed $z$-score exceeds $\tau$. We define overall detectability across all possible thresholds as follows:
[…]
This integral computes the AUROC, which quantifies the ability of the watermarking method to differentiate between human-written and watermarked code. A higher AUROC value (closer to 1) indicates stronger detectability, whereas an AUROC of 0.5 implies that the distributions of human-written and watermarked code are indistinguishable.
4.3 Naturalness
A watermark that significantly alters token distributions can be identified and removed through pre-processing. We adopt perplexity as a direct measure of token-level fluency, which aligns closely with how language models predict code tokens. Other code-style metrics exist, but we rely on perplexity because it has been widely recognized in language modeling research as a reliable indicator of naturalness. A well-designed watermark minimizes perplexity shifts, ensuring that watermarked code remains indistinguishable from non-watermarked code.
For a given set of watermarked code sequences $C_{wm}$, perplexity is computed as:
$\text{PPL}(C_{wm})=\frac{1}{|C_{wm}|}\sum_{j=1}^{|C_{wm}|}\exp\left(-\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\log P(w_{i}^{(j)}|w_{<i}^{(j)})\right),$ (8)
where $N_{j}$ is the number of tokens in the j-th sample of $C_{wm}$, and $P(w_{i}^{(j)}|w_{<i}^{(j)})$ represents the probability of token $w_{i}^{(j)}$ given its preceding context. An effectively concealed watermark minimize perplexity shifts, ensuring imperceptibility. We define the naturalness metric as:
$\mathrm{Naturalness}(C_{wm},C)=1-\frac{|\text{PPL}(C_{wm})-\text{PPL}(C)|}{\text{PPL}(C)},$ (9)
This metric quantifies the perplexity-based difference between LLM-generated code with $C_{wm}$ and without $C$ a watermark. The metric normalizes the absolute perplexity difference by dividing it by the perplexity of non-watermarked code $C$, ensuring invariance. A value of 1 indicates that $C$ and $C_{wm}$ have identical perplexities, implying minimal divergence despite watermark insertion. Conversely, a value close to 0 suggests substantial perplexity divergence, indicating a significant impact of watermark insertion on code generation. By scaling the perplexity difference between 0 and 1, this metric provides a clear and interpretable measure of similarity.
Based on these three criteria, we propose CWEM, a comprehensive evaluation metric for assessing code watermarking effectiveness. Our metric not only enables comparison across different watermarking methods but also offers flexibility through its weighted structure to adjust varying requirements in different application contexts.
5 Results and Analysis
5.1 Experiment Settings
Datasets
For our main experiments, we use the HumanEval+ and MBPP+ datasets (Liu et al., 2023b) for Python code completion and generation tasks, following prior work (Lee et al., 2024). In addition, we incorporate the HumanEvalPack (Muennighoff et al., 2024) to evaluate the performance of our approach across three programming languages: Python, C++, Java. Table 1 shows the data statistics of the datasets we used in our experiments.
Datasets Problems Test Cases Avg Tokens
MBPP+ 399 105 87
HumanEval+ 164 764 57
HumanEvalPack-C++ 164 764 100
HumanEvalPack-Java 164 764 97

### v2

@@ Abstract
Identifying LLM-generated code through watermarking poses a challenge in preserving functional correctness. Previous methods rely on the assumption that watermarking high-entropy tokens effectively maintains output quality. Our analysis reveals a fundamental limitation of this assumption: syntax-critical tokens such as keywords often exhibit the highest entropy, making existing approaches vulnerable to logic corruption. We present STONE, a syntax-aware watermarking method that embeds watermarks only in non-syntactic tokens and preserves code integrity. For its rigorous assessment, we also introduce STEM, a comprehensive framework that balances three critical dimensions: correctness, detectability, and imperceptibility. Across Python, C++, and Java, STONE preserves correctness, sustains strong detectability, and achieves balanced performance with minimal overhead. Our implementation is available at https://anonymous.4open.science/r/STONE-watermarking-AB4B/.
@@
  Jungin Kim   Shinwoo Park   Yo-Sub Han Yonsei University, Seoul, Republic of Korea jungin.kim@yonsei.ac.kr, pshkhh@yonsei.ac.kr, emmous@yonsei.ac.kr
@@ 1 Introduction
The rapid development of LLMs has significantly improved the ability to generate human-like text and code (Nam et al., 2024; Kazemitabaar et al., 2023; Liu et al., 2023a; Sun et al., 2024; Guo et al., 2024a; Wang et al., 2024; Wei et al., 2023; Tang et al., 2023; Guo et al., 2024b). These advancements have unlocked new possibilities across various fields, including software development and automated code generation. However, this progress has raised new challenges in tracing the provenance of generated code (Yang et al., 2023; Li et al., 2023). As LLM-generated code becomes increasingly indistinguishable from human-written code, researchers have explored watermarking techniques to address this issue. As illustrated in Figure 1, existing approaches often modify syntax-critical tokens such as keywords, and this modification introduces syntax errors or alters program behavior. In contrast, STONE preserves these tokens and embeds watermarks only into non-syntactic elements, thereby maintaining structural and functional integrity.
Figure 1: Motivating example of STONE watermarking. Existing methods modify syntax tokens and cause syntax errors, whereas STONE preserves syntax and embeds watermarks without breaking code structure.
@@
LLM watermarking involves embedding specific patterns into the generated output during the generation process to help identify its origin (Sander et al., 2024; Golowich and Moitra, 2024; Hu and Huang, 2024; Giboulot and Furon, 2024; Li et al., 2024b; Pang et al., 2024; Zhou et al., 2024; Fu et al., 2024; Panaitescu-Liess et al., 2025). These patterns remain imperceptible to humans and detectable through algorithmic analysis, thereby promoting greater transparency and accountability in AI-generated content (Guo et al., 2024c; Hou et al., 2024). Among various approaches, the green and red list-based watermarking is the most widely adopted (Kirchenbauer et al., 2023; Zhao et al., 2023; Lee et al., 2024; Lu et al., 2024; Guo et al., 2024c). These approaches divide token candidates into two lists (green and red) and biases the selection process toward tokens in the green list. Detection involves measuring the proportion of green list tokens in the generated output. While effective for natural language tasks, this approach does not easily apply to code generation, where structural and functional correctness are crucial (Lee et al., 2024; Hoang et al., 2024). Changes to syntax-critical tokens can cause compilation errors or alter program behavior.
We propose STONE (Syntax TOkeN preserving codE watermarking), a method that selectively embeds watermarks in non-syntax tokens. By excluding tokens that are critical to code execution, STONE reduces the risk of functional degradation while embedding robust, detectable patterns.
Additionally, we introduce STEM (Summative Test framework for Evaluating code-waterMarking). Existing studies on code watermarking rely on separate evaluation criteria, using different metrics and measurement protocols across methods, hindering fair and consistent comparison of performance. STEM integrates correctness, detectability, and imperceptibility into a unified framework, providing a fair basis for evaluating overall performance while allowing task-specific weighting when needed. We employ STEM to evaluate STONE and compare it with state-of-the-art watermarking methods. Experiments across Python, C++, and Java show that STONE preserves functional correctness, achieves balanced performance across all evaluation criteria, and consistently outperforms prior methods in efficiency.
Figure 2: Entropy values by token type. Tokens that do not fall under keywords, whitespace, types, delimiters, or operators are categorized as etc tokens. The etc category is the primary target for our proposed STONE watermarking method. Token entropy is measured using Qwen2.5-Coder-7B.
2 Motivation and Preliminary Analysis
@@
The recent code watermarking method SWEET (Lee et al., 2024) accounts for structural and syntactic constraints by embedding watermarks only in high-entropy tokens, thereby minimizing their impact on code quality. In our preliminary analysis, we group code tokens into several categories and compare their average entropy values to see if high-entropy tokens have a smaller impact on code quality. We consider two popular benchmarks, MBPP+ and HumanEval+, for the Python code generation and completion. We classify each Python code into one of five categories (keywords, whitespace, types, delimiters, and operators) where any remaining tokens are put into the etc. category. A detailed explanation is provided in Appendix A.
@@
where $H_{t}$ is the entropy value at generation step $t$, $V$ is the vocabulary of the language model, $y_{<t}$ is the sequence of tokens generated before step $t$, and $P(y_{t}=v_{i}\mid y_{<t})$ is the probability assigned by the model to token $v_{i}$ as the next token. The entropy $H_{t}$ measures the uncertainty of the next token prediction. A high entropy value indicates that the model is uncertain and many tokens have comparable probabilities. Conversely, low entropy implies that one token has a dominant probability.
@@
Since keywords have high entropy, embedding watermarks in high-entropy tokens, as in SWEET, carries the risk of placing watermarks on keywords. Keywords are essential to the syntax and logic of the code, thus watermarking them could negatively impact the quality and functionality of the code. In comparison, the etc category, while less critical to the functionality of the code, has relatively high entropy. Therefore, embedding watermarks in etc tokens (i.e., non-syntactic tokens) may be a more suitable approach for preserving code quality.
3 Methodology
3.1 Syntax-Aware Watermarking: STONE
STONE is a code watermarking method specifically designed to preserve the quality of generated code. This section describes how STONE performs watermark insertion and detection. We describe the insertion process, where watermarks are embedded only in non-syntactic tokens while keeping critical elements such as keywords, operators, and delimiters unchanged to preserve functionality and structure. Detection is performed by analyzing the distribution of green tokens among non-syntax elements.
Inserting Watermarks
We define the keywords, whitespace, types, delimiters, and operators as functionally important syntactic elements and embed watermarks only in tokens that do not belong to these categories. Algorithm 1 describes how STONE embeds watermarks during code generation. Given a prompt $\mathbf{x}$ and $\mathbf{y}_{:t}$, the tokens generated up to the previous time step, the language model $f_{\text{LM}}$ computes the initial probability distribution $\mathbf{p}_{t}$ for the token to be generated at the current time step $t$. STONE samples a candidate token $\tilde{y}_{t}$ based on this initial probability distribution. If the sampled token is not in the syntax element set, STONE divides the vocabulary $\mathcal{V}$ into a green/red list according to the green token ratio $\gamma$. It then increases the likelihood of sampling tokens from the green list by adding a constant $\delta$ to their logit values. Using the adjusted logits, STONE updates the probability distribution and samples the final token $y_{t}$ for time step $t$.
Algorithm 1 Insertion Algorithm of STONE
1:Prompt $\mathbf{x}$, previously generated tokens $\mathbf{y}_{:t}=(y_{0},\dots,y_{t-1})$, vocabulary set $\mathcal{V}$, language model $f_{\text{LM}}$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, logit adjustment factor $\delta>0$
@@ 3.1 Syntax-Aware Watermarking: STONE
7:   Compute a hash of token $y_{t-1}$ to use as a seed
8:   Split $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
@@
Detecting Watermarks
We assess how many green tokens are present in the code, and if this value surpasses a certain threshold, we conclude that the code was generated by an LLM. When identifying potential green tokens, we focus only on non-syntactic elements. Algorithm 2 details the watermark detection process. $N^{E}$ represents the number of etc tokens (non-syntax tokens) in the code, and $N^{E}_{G}$ denotes the number of green tokens among them. We replicate the division of the vocabulary into green and red lists at each time step during code generation to determine whether a specific token is a green token. If the $z$-score exceeds the predefined threshold $z_{\text{threshold}}$, the code is determined to be generated by an LLM.
@@
7:   Split $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
@@
9:     $N^{E}_{G}\leftarrow N^{E}_{G}+1$ $\triangleright$ Token matches the green list
@@
3.2 Unified Evaluation Framework: STEM
Prior studies on code watermarking have typically evaluated functional correctness, detectability, and imperceptibility as separate criteria (Lee et al., 2024; Yang et al., 2024; Li et al., 2024a; Guan et al., 2024). However, evaluating each criterion in isolation makes it difficult to capture the trade-offs between them and to identify methods with balanced performance. For example, a watermarking technique may enhance detectability at the expense of correctness, resulting in incomplete or conflicting assessments. This work addresses this limitation by introducing STEM, a balanced evaluation framework that integrates correctness, detectability, and imperceptibility into a unified score. STEM combines the three criteria using weights $\alpha$, $\beta$, and $\zeta$ satisfying $\alpha+\beta+\zeta=1$, enabling direct comparison of methods in terms of overall balance rather than isolated metrics.
$\text{{STEM}}=$ $\;\alpha\cdot\mathrm{Correctness}(C_{wm})$ (2)
$+\beta\cdot\mathrm{Detectability}(C_{wm},C_{H})$
$+\zeta\cdot\mathrm{Imperceptibility}(C_{wm},C).$
Correctness
Maintaining correctness is crucial for preventing watermarking from compromising code functionality. Let $C_{wm}$ denote the set of watermarked code samples. We measure functionality preservation using the unbiased estimator of pass@k Chen et al. (2021):
@@ 3.2 Unified Evaluation Framework: STEM
where $n$ and $c$ denote the total and correct generations per task, respectively. A higher score indicates that watermark insertion does not hinder program execution.
Detectability
Detectability measures how effectively a watermark can be identified in generated code. We assess detectability using a $z$-score that quantifies the proportion of green-list tokens and compute AUROC between human-written and watermarked corpora.
@@
For a generated sequence $X_{wm}=(X_{0},X_{1},\dots,X_{N-1})$, the $z$-score of green-token ratio is defined as
$z(X_{wm})=\frac{|X_{wm}|_{G}-\gamma|X_{wm}|}{\sqrt{\gamma(1-\gamma)|X_{wm}|}},$ (4)
where $|X_{wm}|_{G}$ is the number of green tokens in $X_{wm}$ and $\gamma$ is the expected green-token ratio in non-watermarked code. Higher $z$ values indicate stronger evidence that the code is watermarked.
Given a corpus of human-written code $C_{H}=\{X^{(i)}_{H}\}$ and watermarked code $C_{wm}=\{X^{(j)}_{wm}\}$, we compute the true- and false-positive rates at threshold $\tau$:
$\text{TPR}(\tau)$ $=\frac{\sum_{j=1}^{J}\mathbf{1}\!\bigl[z(X_{wm}^{(j)})>\tau\bigr]}{J}.$ (5)
$\text{FPR}(\tau)$ $=\frac{\sum_{i=1}^{I}\mathbf{1}\!\bigl[z(X_{H}^{(i)})>\tau\bigr]}{I}.$ (6)
The area under the ROC curve (AUROC) is then
$\mathrm{Detectability}(C_{wm},C_{H})=\int_{0}^{1}\text{TPR}(\text{FPR})\,d(\text{FPR}),$ (7)
where higher AUROC values indicate stronger detectability between watermarked and human-written code.
Imperceptibility
Watermarking should preserve token probability distributions consistent with normal model outputs (Huang and Wan, 2025). In code watermarking, Sun et al. (2023) assessed imperceptibility through post-hoc detection combining human and automated evaluations, which does not quantify distributional distortion. Following perplexity-based approaches in text watermarking (Kirchenbauer et al., 2023) and studies including code generation (Magnusson et al., 2024), we define imperceptibility as the deviation in token probability distributions introduced by watermarking, measured through perplexity shifts computed with code LLMs.
Let $C_{wm}$ and $C$ denote the corpora of watermarked and non-watermarked code, respectively. Perplexity is computed as
$\operatorname{PPL}(C_{wm})=\begin{aligned} &\frac{1}{\lvert C_{wm}\rvert}\sum_{j=1}^{\lvert C_{wm}\rvert}\exp\!\bigl(-A_{j}\bigr),\\ &A_{j}=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\log P\!\left(y_{i}^{(j)}\mid y_{<i}^{(j)}\right).\end{aligned}$ (8)
Higher values indicate that watermark insertion produces little distributional deviation from normal code generation.
4 Experimental Setup
Datasets.
For our experiments, we evaluate our approach across three programming languages—Python, C++, and Java—using four benchmark datasets: HumanEval+ and MBPP+ (Liu et al., 2023b) for Python, and HumanEvalPack (Muennighoff et al., 2024) for C++ and Java. Table 1 presents token-length statistics of the evaluation datasets, computed with the Qwen2.5-Coder-7B tokenizer. The datasets span from short snippets to long sequences, enabling assessment of our method across varying complexity.
Dataset Problems Test Cases Max Min Mean Std.
MBPP+ 399 105 341 11 40.25 34.91
HumanEval+ 164 764 558 43 188.28 87.32
HEP-C++ 164 764 697 42 223.10 107.57
HEP-Java 164 764 660 56 237.36 105.09